OpenStudy (anonymous):

Gold will dissolve in the acid mixture known as aqua regia (a mixture of nitric and hydrochloric acids) according to the reaction: Au + HNO3 + 3HCl → AuCl3 + NO + 2H2O. How many grams gold (III) chloride will be produced in this reaction when one starts with 5 mg of gold? (1,000 mg = 1g)

OpenStudy (chmvijay):

can u tell me gold atomic weight

OpenStudy (chmvijay):

and AuCl3 molecular weight

OpenStudy (anonymous):

Au should be 196.96. Cl is 35.45. You can easily find AuCl3

OpenStudy (chmvijay):

u find and tell me for AuCl3

OpenStudy (chmvijay):

quick please

OpenStudy (anonymous):

303.31

OpenStudy (chmvijay):

ok 196.96 g of Au give rise to 303.31 gram of AuCl3 then 5*10-3 g of Au give rise to =303.31*5*10-3 /196.96 =0.0076 g of Aucl3

OpenStudy (chmvijay):

u got it or no

OpenStudy (anonymous):

I got it. Thanks!

OpenStudy (chmvijay):

ur welcome